# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 5.1.3. The minimizer satisfies a Liouville type PDE
assuming that the normalization condition 5.9 holds. The vector $\boldsymbol{\lambda}\in\mathbb{R}^{3},$ depending on $v,$ is uniquely determined by the condition that
@@
To see this, we express
$\delta(v)\coloneqq\delta(\nu)\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\nu),\,\,\,\nu\coloneqq e^{2v}\omega_{\text{FS}}$
It follows from classical convex analysis that $\delta(\nu)$ is differentiable along smooth curves $\nu_{t}$ in $\mathcal{P}(X).$ Indeed, $D(\cdot|\nu)$ is strictly convex and not identically $\infty$ on the non-empty convex and compact subspace $\mathcal{P}(X)_{0}$ of $\mathcal{P}(X),$ which ensures that the inf above is attained at unique minimizer $\mu\in\mathcal{P}(X)_{0}.$ Moreover, when $\nu_{0}:=\nu,$
$\frac{d\delta(\nu_{t})}{dt}_{|t=0}=\left\langle\log(\frac{\mu}{\nu}),\frac{d\nu_{t}}{dt}\right\rangle.$
Furthermore, the minimizer $\mu$ is given by
@@
where $\boldsymbol{\lambda}\in\mathbb{R}^{3}$ is uniquely determined by the condition that $e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu\in\mathcal{P}(X)_{0}.$ Indeed, the existence and uniqueness of $\boldsymbol{\lambda}$ follows directly from the fact that the function $\boldsymbol{\lambda}\mapsto\log\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu$ on $\mathbb{R}^{3}$ is strictly convex and proper (using that $\nu$ and $\omega_{\text{FS }}$are mutually absolutely continuous) and thus admits a critical point $\boldsymbol{\lambda}.$ Moreover, a direct computation reveals that the measure $\mu$ in $\mathcal{P}(X)_{0}$ defined by formula 5.12 is a critical point of $D(\cdot|\nu)$ in $\mathcal{P}(X)_{0}$ and thus minimizers $D(\cdot|\nu).$ Finally, taking $\nu_{t}$ to be the curve corresponding to $v+t\dot{v}$ proves formula 5.10.
@@ 5.1.4. The minimizer $v$ is constant
To conclude the proof it will be enough to show that any minimizer $v$ of $\mathcal{F}_{\gamma}+\delta$ on $W^{2,1}(X)$ is constant. By the uniqueness result for the Liouville equation on $S^{2}$ for $\gamma\in]1,2]$ in [72] it is enough to show that $\boldsymbol{\lambda}=0.$ To this end first recall that, by a result of Kazdan-Warner [39, formula 2.11] (and the normalizations in Lemma 4.1) any solution $u$ of the equation
@@
Indeed, taking $F$ to be a rotation of $S^{2}$ having the property that $F^{*}\boldsymbol{\lambda}\cdot\boldsymbol{x}=\lambda x_{3}$ for some $\lambda\in\mathbb{R}$ we can replace $v$ with $F^{*}v.$ Now, applying the Kazdan-Warner condition 5.13 to $i=3$ gives, by multiplying both sides with $\lambda,$
@@
Indeed, setting $f:=\lambda x_{3}$ we have, by the defining property 5.11 of $\lambda,$ that $\left\langle fe^{f}\right\rangle=0.$ But, by Jensen’s inequality
@@
All in all this means that the bracket in the rhs of formula 5.14 is non-negative and strictly positive unless $\lambda=0.$ Since the same thing goes for the lhs in formula 5.14 and $(2-2\gamma)<0$ we deduce that $\lambda=0,$ as desired.
@@
$\gamma^{-1}F_{\gamma}(\mu)=\mathcal{D}_{\gamma}(u_{\mu})+\gamma^{-1}D(\mu|\mu_{\gamma u_{\mu}}),$
@@
$\gamma^{-1}F_{\gamma}(\mu)\geq\mathcal{D}_{\gamma}(u_{\mu})+\delta_{\gamma}(u_{\mu}).$
@@
Thus, if $\mu$ minimizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ then, by the first part of Theorem 5.1, proved above and the the duality formula 5.1, $u_{\mu}$ minimizes $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ on $W^{2,1}(X).$ But, as shown in Section 5.1.4, this implies that $u_{\mu}$ is constant, i.e. $\mu=\omega_{\text{FS}},$ as desired.

### v3

@@ 5.1.3. The minimizer satisfies a Liouville type PDE
assuming that the normalization condition 5.12 holds. The vector $\boldsymbol{\lambda}\in\mathbb{R}^{3},$ depending on $v,$ is uniquely determined by the condition that
@@
Formula 5.13 follows directly from the following basic lemma applied to $\nu=e^{2v}\omega_{\text{FS}}$ and the curve $\nu_{t}$ corresponding to $v+t\dot{v}.$
Lemma 5.6.
Given $\nu\in\mathcal{P}(X)$ consider the “entropic distance” from $\nu$ to $\mathcal{P}(X)_{0}$ defined by
$\delta(\nu)\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\nu)$
The function $\delta(\nu)$ is differentiable along smooth curves $\nu_{t}$ in $\mathcal{P}(X).$ Morover, the inf defining $\delta(\nu)$ is attained unique $\mu\in\mathcal{P}(X)_{0}$ and $\mu$ is given by
@@
where $\boldsymbol{\lambda}\in\mathbb{R}^{3}$ is uniquely determined by the condition that $e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu\in\mathcal{P}(X)_{0}.$
Proof.
This is, without doubt, essentially well-known, but we provide a simple proof. First, since $D(\cdot|\nu)$ is strictly convex and not identically $\infty$ on the non-empty convex and compact subspace $\mathcal{P}(X)_{0}$ of $\mathcal{P}(X),$ inf above is attained at unique minimizer $\mu\in\mathcal{P}(X)_{0}.$ As a consequence, it follows from general principles (see the appendix in [23]), that for any curve $\nu^{t}$ coinciding with $\nu$ at $t=0,$
$\frac{d\delta(\nu^{t})}{dt}_{|t=0}=\left\langle\log(\frac{\mu}{\nu}),\frac{d\nu^{t}}{dt}\right\rangle.$
Next, the existence and uniqueness of $\boldsymbol{\lambda}$ follows directly from the fact that the function $\boldsymbol{\lambda}\mapsto\log\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu$ on $\mathbb{R}^{3}$ is strictly convex and proper (using that $\nu$ and $\mu$ are mutually absolutely continuous) and thus admits a critical point $\boldsymbol{\lambda}.$ Moreover, a direct computation reveals that the measure $\mu$ in $\mathcal{P}(X)_{0}$ defined by formula 5.15 is a critical point of $D(\cdot|\nu)$ in $\mathcal{P}(X)_{0}$ and thus minimizers $D(\cdot|\nu).$
@@ 5.1.4. The minimizer $v$ is constant
To conclude the proof it will be enough to show that any minimizer $v$ of $\mathcal{F}_{\gamma}+\delta$ on $W^{2,1}(X)$ is constant. By the uniqueness result for the Liouville equation on $S^{2}$ for $\gamma\in]1,2]$ in [80] it is enough to show that $\boldsymbol{\lambda}=0.$ To this end first recall that, by a result of Kazdan-Warner [48, formula 2.11] (and the normalizations in Lemma 4.1) any solution $u$ of the equation
@@
Indeed, taking $F$ to be a rotation of $S^{2}$ having the property that $F^{*}\boldsymbol{\lambda}\cdot\boldsymbol{x}=\lambda x_{3}$ for some $\lambda\in\mathbb{R}$ we can replace $v$ with $F^{*}v.$ Now, applying the Kazdan-Warner condition 5.16 to $i=3$ gives, by multiplying both sides with $\lambda,$
@@
Indeed, setting $f:=\lambda x_{3}$ we have, by the defining property 5.14 of $\lambda,$ that $\left\langle fe^{f}\right\rangle=0.$ But, by Jensen’s inequality
@@
All in all this means that the bracket in the rhs of formula 5.17 is non-negative and strictly positive unless $\lambda=0.$ Since the same thing goes for the lhs in formula 5.17 and $(2-2\gamma)<0$ we deduce that $\lambda=0,$ as desired.
@@
$\gamma^{-1}F_{-\gamma}(\mu)=\mathcal{D}_{\gamma}(u_{\mu})+\gamma^{-1}D(\mu|\mu_{\gamma u_{\mu}}),$
@@
(5.18) $\gamma^{-1}F_{-\gamma}(\mu)\geq\mathcal{D}_{\gamma}(u_{\mu})+\delta_{\gamma}(u_{\mu}).$
@@
Thus, if $\mu$ minimizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ then, by the first part of Theorem 5.1, proved above and the the duality formula 5.4, $u_{\mu}$ minimizes $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ on $W^{2,1}(X).$ But, as shown in Section 5.1.4, this implies that $u_{\mu}$ is constant, i.e. $\mu=\omega_{\text{FS}},$ as desired.
5.2. Application to optimal stability results for the sharp log HLS inequality on $S^{2}$
Starting with the seminal stability result [32] for the sharp Sobolev inequality in $\mathbb{R}^{n},$ for $n\geq 3,$ a wide variety of quantitative stability results for various sharp functional inequalities $\text{LHS }(f)\leq\text{RHS }(f)$ have been established of the form
(5.19) $\text{RHS }(f)-\text{LHS }(f)\geq\kappa d(f,M),\,\,\,\,d(f,M)\coloneqq\inf_{g\in M}d(f,g)$
where $M$ denotes the set of optimizers, i.e. all $f$ for which $\text{LHS }(f)=\text{RHS }(f),$ $\kappa$ is a positive constant, that we shall call the stability constant and $d(f,g)$ is a given divergence, i.e. $d(f,g)\geq 0$ with equality iff $f=g.$ In the original case of the Sobolev inequality [32], the divergence was the squared Dirichlet-norm, but $\kappa$ was non-explicit. This stability result was extended in [50] to all sharp fractional Sobolev inequalities in $\mathbb{R}^{n}$ of order $\alpha\in]0,n/2[$ (coinciding with the ordinary Sobolev inequality when $n\geq 3$ and $\alpha=1),$ with a (non-explicit) constant $\kappa_{\mathcal{S}_{n,\alpha}}$ and then to the corresponding dual inequalities, the Hardy–Littlewood–Sobolev inequalities (HLS) of order $\alpha\in]0,n/2[$ in [46] using duality:
(5.20) $C_{n,\alpha}^{*}\left\|f\right\|_{L^{\frac{2n}{n+2\alpha}}(\mathbb{R}^{n})}^{2}-\left\|(-\Delta)^{-\alpha/2}f\right\|_{(\mathbb{R}^{n})}^{2}\geq\kappa_{\mathcal{S}_{n,\alpha}^{*}}\left\|f-M_{S_{n,\alpha}^{*}}\right\|_{L^{\frac{2n}{n+2\alpha}}(\mathbb{R}^{n})}^{2},$
where $C_{n,\alpha}^{*}$ denotes the sharp constant in the HLS inequality (which is explicitly known), $M_{S_{n,\alpha}^{*}}$ is the corresponding $(n+2)$-dimensional submanifold of optimizers and $\kappa_{\mathcal{S^{*}}_{n,\alpha}}$ is explicitly expressed in terms of $\kappa_{S_{n,\alpha}}.$ Since, as shown in [44], differentiating the HLS inequalities at the critical value $\alpha=n/2$ (in an appropriate one-sided sense) yields the log HLS inequality it is, as pointed out in [46, Section 3.2], tempting to try to extract a stability result for the log HLS inequality from the stability result 5.20. However, this would require detailed information about the dependence of $\kappa_{S_{n,\alpha}}$ on $\alpha.$ In the case $n=2$ (so that $\alpha=1$ is the critical case) a quantitative stability result for the log HLS inequality was first established in [45] and recently considerably improved in [47], showing that, after stereographic projection to the two-sphere $S^{2}$ and using our notation,
(5.21) $D(\mu)-E(\mu)\geq\frac{1}{8}\left\|\mu-M\right\|_{\text{TV}}^{2},\,\,\,\mu\in\mathcal{P}(S^{2}),\,\,D(\mu)<\infty,$
where $M$ denotes the 33-dimensional submanifold of $\mathcal{P}(S^{2})$ consisting of all optimizers $\mu$ and $\left\|\mu-\nu\right\|_{\text{TV}}^{2}$ is the total variation norm. Since the entropy $D(\mu)$ arises when differentiating the $L^{\frac{2n}{n+2\alpha}}$-norm at the critical case $\alpha=n/2,$ it seems natural to ask if the the norm $\left\|\mu-\nu\right\|_{\text{TV}}^{2}$ can be replaced by the relative entropy $D(\mu,\nu),$ i.e. by the Kullback–Leibler divergence. In fact, as next shown, this follows from Theorem 5.1 with an optimal stability constant $\kappa$:
Corollary 5.7.
Denote by $\Delta$ the Laplacian wrt any Riemannian metric on $S^{2}$ that is conformally equivalent to the one induced by the Euclidean metric on $\mathbb{R}^{3}$ and by $dV$ the corresponding volume form. The following stability inequalities hold for any $\mu\in\mathcal{P}(S^{2}),$ such that $D(\mu)<\infty,$
(5.22) $D(\mu)-E(\mu)\geq 1\cdot\left\|(-\Delta)^{1/2}(\mu-M)\right\|_{L^{2}(S^{2},dV_{g})}^{2},$
and
(5.23) $D(\mu)-E(\mu)\geq\frac{1}{2}D\left(\mu|M\right),$
Moreover, the constants $1$ and $1/2$ are optimal, i.e. the stability inequalities above fail when these constants are replaced by any strictly larger constants.
Proof.
By Theorem 5.1
$D(\mu)-E(\mu)\geq 1\cdot\inf_{g\in\mathcal{G}}E(g^{*}\mu),$
where $1$ is the optimal constant. Moreover, by formula 4.3, $E(g^{*}\mu)=\left\|(g^{*}\mu-\mu_{0})\right\|_{H^{-1}}^{2},$ which equals $\left\|(\mu-g_{*}\mu_{0})\right\|_{H^{-1}}^{2},$ since the $H^{-1}$-norm is $\mathcal{G}$-invariant. Equivalently, $E(g^{*}\mu)=\left\|(-\Delta)^{1/2}(\mu-g_{*}\mu_{0})\right\|_{L^{2}(S^{2},dV)}^{2}.$ Since $M$ coincides with the $\mathcal{G-}$orbit of $\mu_{0}$ this proves the first stability inequality 5.22 and shows that $1$ is the optimal stability constant.
To prove the second stability inequality 5.23 first assume that $\mu\in\mathcal{P}(X)_{0}.$ Then Theorem 5.1 for $\gamma=2$ yields the first inequality in the following equivalence
$F_{-2}(\mu)\coloneqq-2E(\mu)+D(\mu)\geq 0\iff F_{-1}(\mu)\coloneqq D(\mu)-E(\mu)\geq\frac{1}{2}D(\mu).$
Now take any $\mu$ in $\mathcal{P}(X)$ such that $D(\mu)<\infty.$ In particular, $E(\mu)<\infty.$ Hence, by Prop 3.4, there exists $g\in\mathcal{G}$ such that $g_{*}\mu\in\mathcal{P}(X)_{0}.$ Moreover, by Prop 3.4 $g_{*}\mu$ minimizes $E(\mu)$ on the $\mathcal{G}$-orbit of $\mu.$ But this means, equivalently, that $g_{*}\mu$ minimizes $D(\mu)$ on the $\mathcal{G}$-orbit of $\mu,$ since $D(\mu)-E(\mu)(\coloneqq F_{-1}(\mu))$ is $\mathcal{G}$-invariant. Hence,
$F_{-1}(\mu)=F_{-1}(g_{*}\mu)\geq\frac{1}{2}D(g_{*}\mu)=\frac{1}{2}\inf_{g\in\mathcal{G}}D(g_{*}\mu)$
Next observe that
(5.24) $\inf_{g\in\mathcal{G}}D(g_{*}\mu)=D(\mu|M).$
Indeed, $D(g_{*}\mu)\coloneqq D(g_{*}\mu|\mu_{0})=D(\mu|g^{*}\mu_{0})$ and $F_{-1}(\mu)=0$ iff there exists $g\in\mathcal{G}$ such that $\mu=g^{*}\mu_{0}.$ Combining the previous two inequalities thus yields the inequality in the corollary. Finally, to prove that the optimality of the stability constant $1/2$ assume, to get a contradiction, that there exists a small positive number $\epsilon$ such that
$F_{-1}(\mu)\geq\frac{1}{2-\epsilon}D(\mu|M).$
By the previous argument this means, equivalently, that $D(\mu)-E(\mu)\geq\frac{1}{2-\epsilon}D(\mu)$ on $\mathcal{P}(X)_{0}$ i.e. that $D(\mu)-\gamma E(\mu)\geq 0$ on $\mathcal{P}(X)_{0}$ for $\gamma\coloneqq(2-\epsilon)/(1-\epsilon)>2,$ which contradicts Theorem 3.13 saying, in particular, that $F_{-\gamma}(\mu)$ is unbounded from below on $\mathcal{P}(X)_{0}$ when $\gamma>2.$
5.2.1. Comparison with the stability inequality in [47] and stability of the sharp Moser-Trudinger inequality
Combining the previous corollary with the sharp Pinsker inequality,
$D(\mu|\nu)\geq\left\|\mu-\nu\right\|_{\text{TV}}^{2}/2,$
yields
(5.25) $D(\mu)-E(\mu)\geq\frac{1}{4}\left\|\mu-M\right\|_{\text{TV}}^{2},$
which thus improves on the stability result 5.21 from [47]. The main reason for this improvement is that the proof of the latter stability in [47] is based on the sharp improvement of Aubin’s constrained Moser–Trudinger inequality established in [80] (recalled in Theorem 4.5), while the stability result in Cor 5.7 is a consequence of the sharp constrained log HLS inequality in Theorem 5.1, which is, a priori, stronger than Theorem 4.5, as discussed in Section 4.0.1. For comparison, let us summarize the proof of 5.21 in [47]. The starting point is Theorem 4.5 from [80] that, as shown in [47, Thm 4.3], implies an explicit stability inequality for the sharp Moser-Trudinger inequality, that in our notation may be expressed as
(5.26) $\mathcal{D}_{1}(u)\geq\frac{1}{2}\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g),\,\,\,\forall u\in W^{2,1}(S^{2})$
(the proof is similar to the proof of Prop 4.6). The stability inequality 5.21 is then deduced, using the general duality theory for stability inequalities developed in [46], combined with a strong Young inequality for the relative entropy established in [47, Thm 3.2]. But here we note that, starting from the stability inequality 5.26, there is, in fact, an alternative argument that yields a stronger stability inequality for the log HLS inequality than the inequality 5.21, but still weaker than the stability inequality 5.25. Indeed, on $W^{2,1}(S^{2})\cap\mathcal{H}(X,\omega_{0})$ the inequality 5.26 is equivalent to the inequality in Prop 4.6, which, as shown in the proof of Prop 4.7, implies that $F_{-3/2}(\mu)\geq 0$ on $\mathcal{P}(X)_{0}.$ Thus implies, precisely, as in the proof of Cor 5.7, that
$D(\mu)-E(\mu)\geq\frac{1}{3}D\left(\mu|M\right),$
giving $D(\mu)-E(\mu)\geq 6^{-1}\left\|\mu-M\right\|_{\text{TV}}^{2},$ by the sharp Pinsker inequality.
Finally, we note that (as in the formulation of [47, Thm 4.3]), the inequality 5.26 is a stability inequality for the sharp Moser–Trudinger inequality of the general form 5.19 on $W^{2,1}(S^{2})/\mathbb{R}$ with explicit stability constant $\kappa=1/8$ wrt the divergence $d(u,v):=\left\|u-v\right\|_{H^{-1}}^{2}.$ It seems thus natural to ask what the value of the corresponding optimal stability $\kappa$ constant is? In this direction it follows from Theorem 1.3 together with the argument in the proof of Prop 4.7, that $\kappa\leq 1/4.$ If the inf over $\mathcal{J}(u\cdot g)$ over $\mathcal{G}$ were attained for $u_{g}:=u\cdot g$ satisfying Aubin’s constraint, then one could conclude from Theorem 4.5 that $\kappa=1/8$ is optimal, using that $\mathcal{D}_{1}$ is $\mathcal{G}$-invariant. However, this is not the case, in general. For example, when $u\in\mathcal{H}(S^{2},\omega_{0})$ the inf in question is, as explained in Section 4.0.1, attained for $u_{g}$ satisfying Tian’s orthogonality constraint, which, in this case is equivalent to $\omega_{u_{g}}\in\mathcal{P}(S^{2})_{0}.$ In contrast, Aubin’s constraint says, in our notation, that $e^{-u_{g}}\omega_{0}\in\mathcal{P}(S^{2})_{0}.$
5.2.2. Comparison with other recent stability results
To the best of our knowledge Cor 5.7 provides the first instance of stability results of the form 5.19, where the optimal stability constant $\kappa$ is explicitly computed, in the context of fractional Sobolev and HLS inequalities or their logarithmic versions. The original compactness argument employed in [32], generalized to fractional Sobolev inequalities (with $\alpha\in]0,n/2[$) in [50], did not yield any explicit value for the corresponding optimal stability constants $\kappa_{S_{n,\alpha}},$ but it did yield an upper bound, showing that $\kappa_{S_{n,\alpha}}\leq 4\alpha/(n+2+2\alpha).$ Only recently an explicit lower bound on the optimal constant $\kappa_{\mathcal{S}_{n,1}}$ (for $n\geq 3)$ was established in [65] of the form $\kappa_{S_{n,1}}\geq\beta/n$ for an explicit constant $\beta,$ whose dimension dependence is thus of the same asymptotic order as the upper bound. By letting $n\rightarrow\infty$ a stability result for the sharp Gaussian log Sobolev inequality in $\mathbb{R}^{n}$ was deduced in [65], with an explicit (but not optimal) constant $\kappa,$ where the corresponding divergence $d$ is the squared Gaussian $L^{2}$-norm. Very recently the explicit lower bound on $\kappa_{\mathcal{S}_{n,1}}$ in [65] (for $n\geq 3)$ was extended to $\kappa_{\mathcal{S}_{n,\alpha}}$ and $\kappa_{\mathcal{S^{*}}_{n,\alpha}}$ in [49] for any $\alpha\in]0,n/2[,$ replacing the constant $\beta$ with a constant $\beta_{\alpha}$ depending on $\alpha.$
The relative entropy $D(\cdot|\cdot)$ (or more precisely, functions of $D(\cdot|\cdot)$) has previously appeared as a divergence in stability results of the form 5.19. For example, under a moment constraint, a stability results for the sharp Gaussian log Sobolev inequality in $\mathbb{R}^{n}$ was established in [66], where the divergence is an explicit function $\Phi(D(\cdot|\cdot))$ of the relative entropy $D(\cdot|\cdot)$ (see also the recent survey [37]).